A die was rolled 600 times with the following results. Use the Chi-square Goodness of Fit Test to test whether the die is balanced. Use a significance level of 0.05. Give the null and alternative hypotheses, test statistic value, decision rule, and p-value.

value

1 2 3 4 5 6

Frequency

87 96 108 89 122 98
Answer #1

Solution:

Given: A die was rolled 600 times with the following results.

We have to use the Chi-square Goodness of Fit Test to test whether the die is balanced.

Level of significance = 0.05

Step 1) State the null and alternative hypotheses.

H0: the die is balanced

Vs

H1: the die is not balanced

Step 2) test statistic value:

Chi square test statistic for goodness of fit

Where

Oi = Observed Counts

Ei =Expected Counts = N / k = 600 / 6 = 100

Outcomes Oi: Observed frequency Ei: Expected Frequency Oi2/Ei
1 87 100 75.69
2 96 100 92.16
3 108 100 116.64
4 89 100 79.21
5 122 100 148.84
6 98 100 96.04
N = 600

Thus

Step 3) Decision rule

Reject null hypothesis H0, if P-value < 0.05 Level of significance , otherwise we fail to reject H0.

Step 4) P-value:

df = k - 1 = 6 -1 = 5

Look in Chi-square table for df = 5 row and find an interval in which fall, then corresponding right tail area is P-value.

fall between 1.610 and 9.236

corresponding right tail area is between 0.100 and 0.900

thus 0.100 < P-value < 0.900

or P-value > 0.100

Since P-value > 0.100 > 0.05 level of significance, we reject H0.

Thus there is not sufficient evidence to conclude that: the die is balanced.